# Group cohesion and passive dynamics of a pair of inertial swimmers with three-dimensional hydrodynamic interactions

### v1

@@ I Introduction
The hydrodynamic (and other) benefits of schooling can only be realized if the appropriate group structure is maintained. Thus, the dynamics that dictate the trajectories of freely moving swimmers become important. These dynamics are governed by the active decision-making and social interactions of swimmers [29, 30], as well as by passive hydrodynamic interactions. A substantial and important body of work on active decision-making and social interactions exists, and it has produced a variety of agent-based models for the interactions between individuals in a group [31, 32, 33, 34]. These models range in complexity, but they tend to neglect the role of hydrodynamic interactions, which are expected to have an outsized effect on the collective motion of swimmers [35]. Here, we focus on passive hydrodynamic interactions. Considering only the hydrodynamic interactions allows us to answer whether a cohesive group can form passively, or whether feedback control is required to maintain a cohesive group. Additionally, a deeper understanding of the hydrodynamic interactions may change the inferences that one makes about social interactions between animals based on observations of groups of animals, leading to improved agent-based models [36], and could also lead to new physics-based models of underwater robots that could be used for improved model-based control of groups.
Lighthill raised the possibility that the order observed within groups in nature may passively arise due to flow-mediated interactions between group members [37]. Several studies have sought to understand the fluid-mediated dynamics of multiple swimmers interacting with each other [38, 39, 40, 36, 41, 42, 43, 44, 17, 20, 45], with nearly all of them being two-dimensional in nature. Ristroph’s work has consistently shown that swimmers in schools passively and stably arrange themselves so that their trajectory has a particular phase offset from incoming vortical wakes [42, 43, 44]. It is important to note that the experiments leading to this conclusion were highly constrained, each model swimmer having only one unconstrained degree of freedom. Additional unconstrained degrees of freedom may qualitatively change the dynamics.
@@
It is unknown how three-dimensional hydrodynamic interactions, as well as the additional degrees of freedom of a three-dimensional world, affect the stability and cohesion of schools. However, it is immediately clear that the three-dimensional dynamics will be different from the two-dimensional dynamics since the velocity field of a three-dimensional swimmer decays with distance $r$ as $r^{-3}$, while the velocity field of a two-dimensional swimmer decays as $r^{-2}$. Any potential differences are important since our current understanding of the role of hydrodynamics in collective motion is based on the two-dimensional scenario.
In this work, we take the first step towards understanding the passive group dynamics that develop under three-dimensional hydrodynamic interactions between swimmers. This understanding is enabled by the development of a simple model of a freely moving inertial swimmer that captures the correct leading-order hydrodynamic interactions between swimmers. We consider the elemental interactions between a pair of swimmers. By focusing on their planar dynamics, we are able to conduct a comparative analysis between two-dimensional and three-dimensional swimmers, thereby revealing the significance of three-dimensionality in schooling behavior. Additionally, we make observations on how the geometry of a swimmer affects the dynamics by comparing swimmers that are long and thin to those that are short and fat.
@@ II.1 Flow induced by one swimmer
The flow induced by the motion of a body can be expressed in a multipole expansion. For a body of constant volume, the leading-order term is a dipole [46, Ch. 4.7], otherwise called a source doublet [47, Ch. 6.4]. This motivates modeling a swimmer as a source-sink pair of equal strength separated by a fixed distance $\ell$, which yields a dipole flow at large distances while also endowing a length scale absent from a doublet. Effectively, our model represents the far field of a swimmer. As will become clear later, the separation between the source and sink unlocks non-trivial rotational dynamics in our model swimmers. The streamlines in a plane containing the source-sink pair are shown in Figure 1. From a physical standpoint, the source represents the fore of a swimmer, which pushes fluid away as the swimmer moves forward. Conversely, the sink represents the rear of a swimmer, where fluid must rush in to fill the void left by the rear as the swimmer moves forward.
Figure 1: Streamlines in a plane containing the swimmer. The swimmer is oriented along the $z$-axis. The source is highlighted in red and the sink in blue.
To validate the source-sink pair as an effective representation of a swimmer, we compare the flow field generated by our model to that generated by a tuna, as calculated in the simulations of Zhu et al. [48]. In their work, Zhu et al. employed a three-dimensional panel method to simulate both fish and fish-like robots in an inviscid flow, which agreed reasonably with experimental measurements. The streamlines produced by the source-sink pair are shown in three different planes in Figures 1 and 2. The qualitative agreement with the streamlines of the higher-fidelity tuna simulations is excellent outside of the wake region (cf. Figure 5 in [48]). Even very close to the body (within a small fraction of the length of the body) the qualitative agreement is excellent. The streamlines in the wake region differ, which is to be expected since our model does not include vortex shedding. Despite this limitation, our approach should reasonably model the far-field hydrodynamic interactions between swimmers. We return to a discussion of the limitations of the model later.
Figure 2: (a)  Cross-sectional view of the streamlines in the plane $z=0.5$. (b)  Cross-sectional view of the streamlines in the plane $z=-0.5$. The swimmer is the same as in Figure 1.
@@
We explored modeling a swimmer as a thin three-dimensional vortex ring. In our model, the ring was discretized into a set of evenly spaced material points, induced velocities were evaluated using the Biot-Savart law as described in [49], the cut-off method was employed to de-singularize induced velocity calculations [50, Ch. 11.1], and we developed a numerical procedure to enforce that the shape of the vortex ring not change in time. Besides being more complex and computationally expensive than the source-sink pair model, we believe that the vortex ring model is less representative of typical swimmers for reasons that will be clarified in the ensuing section. Thus, the work described here focuses on the source-sink pair model.
@@ II.2 Dynamics of freely moving swimmers
At a displacement $\mathbf{r}$ away from a source, the velocity induced by it is:
@@
When more than one swimmer is present, they will mutually affect each other’s motions. For each swimmer, we take the velocity of the source to be equal to the sum of the self-propelled velocity and the velocity induced by the sources and sinks of all other swimmers. The velocity of the sink is calculated the same way. In other words, the source and sink are effectively treated as material points. Since swimmers have approximately the same density as the surrounding fluid, treating the velocities of the source and sink in this way is reasonable. In contrast, this may be a poor approach for flying animals (depending on their size) since they are typically significantly denser than the air they fly in; the relatively high inertia of a flying animal would lead it to respond much more slowly to its surrounding flow.
@@
To find $\lambda$, start with $\mathbf{x}_{f}-\mathbf{x}_{b}=\ell\mathbf{n}$ and take the time derivative of this equation to arrive at
@@
Figure 5: (a)  A gradient along the body of the crossflow results in rotation. (b)  A gradient across the body of the body-parallel velocity does not result in rotation.
Figure 6: (a)  A gradient along the body of the crossflow does not result in rotation. (b)  A gradient across the body of the body-parallel velocity results in rotation.
@@
III Planar dynamics of a pair
We present numerical examples of the three-dimensional dynamic interactions between a pair of identical swimmers. Such pairwise interactions can be thought of as the elemental interactions in a large school. To facilitate comparison with two-dimensional models, we consider planar dynamics, where the swimmers’ bodies are co-planar initially and, therefore, for all time. This also allows us to fully characterize the phase space. In what follows, lengths have been non-dimensionalized by the swimmer’s length $\ell$ and velocities have been non-dimensionalized by the self-propelled speed $U$.
@@ III.1 Symmetry reduction
With two swimmers, the dimension of the state space is 12, with six variables for each swimmer (three for the position of its center, and three for the unit normal describing its orientation). For the planar motion of two swimmers, however, leveraging symmetry allows us to describe the motion of the pair using only three variables. Specifically, due to a translational symmetry, only the separation between the swimmers’ centers (rather than their absolute positions) affects the dynamics. Similarly, a rotational symmetry makes it so that only the relative angle between the swimmers (rather than their absolute orientations) affects the dynamics. Without loss of generality, we place the swimmers in the $x$-$y$ plane, with one swimmer initially oriented in the positive $y$-direction, separated from each other by $(\Delta x,\Delta y)$, and with a relative angle $\Delta\theta$ between them, as in Figure 7. In this smaller three-dimensional symmetry-reduced phase space, we can fully characterize the possible dynamics.
@@
III.2 Three-dimensional model
@@
Figure 9: Planar trajectories of two swimmers for different initial conditions with $\Delta\theta=0$. (a)  $\Delta\theta=0$, $\Delta x=0.5$, $\Delta y=0$. In (a), the two swimmer’s bodies are represented by arrows whose colors represent time (darker is a later time). (b)  $\Delta\theta=0$, $\Delta x=5$, $\Delta y=0.5$. (c)  $\Delta\theta=0$, $\Delta x=1$, $\Delta y=1$. (d)  $\Delta\theta=0$, $\Delta x=0.5$, $\Delta y=1$.
@@
III.3 Two-dimensional model
@@
III.4 Geometric effects
@@
Finally, we briefly study the effects of geometry by comparing long and thin swimmers (source-sink pairs) to short and fat swimmers (vortex ring model). The phase diagram for the planar dynamics of two initially aligned ($\Delta\theta=0$) vortex-ring swimmers is shown in Figure 15. The region where the vortex rings overlap is excluded. We observe four types of motions: in-line motion, divergence, collision, and oscillatory motion. Oscillatory motion is qualitatively similar to the braid-like motion in Figure 9, except that the paths of the swimmers do not intersect, instead following an alternating pattern of convergence and divergence. Although the types of motions are similar to what is observed for the source-sink model, the phase diagram differs substantially. First, the set of configurations leading to a cohesive group is much smaller than for the source-sink model. Second, cohesive configurations have small vertical separations, while for the source-sink model they had large vertical separations. Third, the vortex rings collide only when $\Delta y=0$; when they collide, they do so by turning head-on towards each other. These qualitative differences in the dynamics of short and fat swimmers compared to long and thin swimmers can be rationalized by appealing to how these two types of swimmers sample the velocity gradient, as was explained in Figures 5 and 6. Interestingly, the dynamics we observe in our vortex ring model are qualitatively similar to the dynamics observed for toroidal swimmers in Stokes flow [51], despite being at opposite ends of the Reynolds number spectrum.

### v2

@@ I Introduction
The hydrodynamic (and other) benefits of schooling can only be realized if the appropriate group structure is maintained. Thus, the dynamics that dictate the motions of freely moving swimmers become important. These dynamics are governed by the active decision-making and social interactions of swimmers [29, 30], as well as by passive hydrodynamic interactions. A substantial and important body of work on active decision-making and social interactions exists, and it has produced a variety of agent-based models for the interactions between individuals in a group [31, 32, 33, 34]. These models range in complexity, but they tend to neglect the role of hydrodynamic interactions, which are expected to have an outsized effect on the collective motion of swimmers [35]. Here, we focus on passive hydrodynamic interactions. Considering only the hydrodynamic interactions allows us to answer whether a cohesive group can form passively, or whether feedback control is required to maintain a cohesive group. Additionally, a deeper understanding of the hydrodynamic interactions may change the inferences that one makes about social interactions between animals based on observations of groups of animals, leading to improved agent-based models [36], and could also lead to new physics-based models of underwater robots that could be used for improved model-based control of groups.
Lighthill famously raised the possibility that the order observed within groups in nature may passively arise due to flow-mediated interactions between group members [37]. Kelly’s earlier analysis of two spheres moving with equal and constant velocities in potential flow lent some credence to this possibility as it showed that the spheres can experience attractive forces, depending on their configuration [38]. Several studies have sought to understand the fluid-mediated dynamics of multiple swimmers interacting with each other [39, 40, 41, 36, 42, 43, 44, 45, 17, 20, 46], with nearly all of them being two-dimensional in nature. Ristroph’s work has consistently shown that swimmers in schools passively and stably arrange themselves so that their motion has a particular phase offset from incoming vortical wakes [43, 44, 45]. It is important to note that the experiments leading to this conclusion were highly constrained, with each model swimmer having only one unconstrained degree of freedom. Additional unconstrained degrees of freedom may qualitatively change the dynamics, as shown in [47, 48, 49], where the two translational degrees of freedom were unconstrained. We expect that unconstrained rotational motion will lead to further changes in the dynamics.
@@
It is unknown how three-dimensional hydrodynamic interactions, as well as the additional degrees of freedom of a three-dimensional world, affect the stability and cohesion of schools. However, it is immediately clear that the three-dimensional dynamics will be different from the two-dimensional dynamics since the velocity field of a three-dimensional swimmer decays with distance $r$ as $r^{-3}$, while the velocity field of a two-dimensional swimmer decays as $r^{-2}$. Any potential differences are important since our current understanding of the role of hydrodynamics in collective motion is based almost solely on the two-dimensional scenario.
In this work, we take the first step towards understanding the passive group dynamics that develop under three-dimensional hydrodynamic interactions between swimmers. This understanding is enabled by the development of a simple model of a freely moving inertial swimmer that captures the correct leading-order hydrodynamic interactions between swimmers. We consider the elemental interactions between a pair of swimmers. By focusing on their planar dynamics, we are able to conduct a comparative analysis between two-dimensional and three-dimensional swimmers, thereby revealing the significance of three-dimensionality in schooling behavior. Symmetry reduction allows us to fully characterize the possible planar dynamics. Additionally, we demonstrate that the non-planar dynamics of a pair are of the same character as the planar dynamics, and that it is possible for a group to be passively cohesive in three dimensions. Finally, we make observations on how the geometry of a swimmer affects the dynamics by comparing swimmers that are long and thin to those that are short and fat.
@@ II.1 Flow induced by one swimmer
The flow induced by the motion of a body can be expressed in a multipole expansion. For a body of constant volume, the leading-order term is a dipole [50, Ch. 4.7], otherwise called a source doublet [51, Ch. 6.4]. This motivates modeling a swimmer as a source-sink pair of equal strength separated by a fixed distance $\ell$, which yields a dipole flow at large distances while also endowing a length scale absent from a doublet. Effectively, our model represents the far field of a swimmer. As will become clear later, the separation between the source and sink unlocks non-trivial rotational dynamics in our model swimmers. The streamlines in a plane containing the source-sink pair are shown in Figure 1. From a physical standpoint, the source represents the fore of a swimmer, which pushes fluid away as the swimmer moves forward. Conversely, the sink represents the rear of a swimmer, where fluid must rush in to fill the void left by the rear as the swimmer moves forward.
Figure 1: (a) Streamlines in a plane containing the swimmer. The swimmer is oriented along the $z$-axis. The source is highlighted in red and the sink in blue. (b)  Cross-sectional view of the streamlines in the plane $z=0.5$. (c)  Cross-sectional view of the streamlines in the plane $z=-0.5$.
To validate the source-sink pair as an effective representation of a swimmer, we compare the flow field generated by our model to that generated by a tuna, as calculated in the simulations of Zhu et al. [52]. In their work, Zhu et al. employed a three-dimensional panel method to simulate both fish and fish-like robots in an inviscid flow, which agreed reasonably with experimental measurements. The streamlines produced by the source-sink pair are shown in three different planes in Figure 1. The qualitative agreement with the streamlines of the higher-fidelity tuna simulations is excellent outside of the wake region (cf. Figure 5 in [52]). Even very close to the body (within a small fraction of the length of the body) the qualitative agreement is excellent. The streamlines in the wake region differ, which is to be expected since our model does not include vortex shedding. Despite this limitation, our approach should reasonably model the far-field hydrodynamic interactions between swimmers. We return to a discussion of the limitations of the model later.
@@
We explored modeling a swimmer as a thin three-dimensional vortex ring. In our model, the ring was discretized into a set of evenly spaced material points, induced velocities were evaluated using the Biot-Savart law as described in [53], the cut-off method was employed to de-singularize induced velocity calculations [54, Ch. 11.1], and we developed a numerical procedure to enforce that the shape of the vortex ring not change in time. Besides being more complex and computationally expensive than the source-sink pair model, we believe that the vortex ring model is less representative of typical swimmers for reasons that will be clarified in the ensuing section. Thus, the work described here focuses on the source-sink pair model.
@@ II.2 Dynamics of freely moving swimmers
At a displacement $\mathbf{r}$ away from a source, the velocity induced by it is
@@
When more than one swimmer is present, they will mutually affect each other’s motions. To determine the hydrodynamic interactions, we treat each swimmer as a dumbbell comprising two beads connected by a rigid rod, the beads respectively co-located with the source and sink. This dumbbell representation is common in the microswimmer literature [55, 56, 57, 58, 59]. In this representation, all the hydrodynamic forces are concentrated on the beads, and the rigid rod serves to keep the length of the dumbbell fixed.
Consider one bead of a swimmer. Let $\mathbf{R}(t)$ be its position, $\mathbf{V}(t)$ be its velocity, and $\mathbf{u}_{0}(\mathbf{x},t)$ be the velocity field induced by the other bead and swimmers in the absence of the bead under consideration. For an incompressible, irrotational, and inviscid flow, the force on the bead is
$\mathbf{F}=\rho\mathcal{V}\left[(1+C_{M})\frac{D\mathbf{u}_{0}}{Dt}\Big{|}_{\mathbf{x}=\mathbf{R}}-C_{M}\frac{\text{d}\mathbf{V}}{\text{d}t}\right],$ (2)
where $\rho$ is the density of the fluid, $\mathcal{V}$ is the volume of the bead, $C_{M}$ is the added mass coefficient of the bead, $\frac{D}{Dt}$ is the material derivative, and $\frac{\text{d}}{\text{d}t}$ is the time derivative following the bead [60].
Applying Newton’s second law to the bead gives
$\frac{\text{d}\mathbf{V}}{\text{d}t}=\frac{1+C_{M}}{1+C_{M}+\Delta\rho/\rho}\frac{D\mathbf{u}_{0}}{Dt}\Big{|}_{\mathbf{x}=\mathbf{R}},$ (3)
where $\Delta\rho=\rho_{s}-\rho$ is the difference between the swimmer’s density $\rho_{s}$ and the fluid’s density. When $\Delta\rho=0$,
$\frac{\text{d}\mathbf{V}}{\text{d}t}=\frac{D\mathbf{u}_{0}}{Dt}\Big{|}_{\mathbf{x}=\mathbf{R}}.$ (4)
That is, the bead moves as a material point if the density of the swimmer is equal to that of the fluid.
For each swimmer, therefore, we take the velocity of the source to be equal to the sum of the self-propelled velocity and the velocity induced by the sources and sinks of all other swimmers. The velocity of the sink is calculated the same way. Since swimmers have approximately the same density as the surrounding fluid, treating the velocities of the source and sink in this way is reasonable. In contrast, this may be a poor approach for flying animals (depending on their size) since they are typically significantly denser than the air they fly in; the relatively high inertia of a flying animal would lead it to respond much more slowly to its surrounding flow.
@@
To find $\lambda$, take the time derivative of $\mathbf{x}_{f}-\mathbf{x}_{b}=\ell\mathbf{n}$ to arrive at
@@
Figure 4: Rotational dynamics of the source-sink pair model. (a)  A gradient along the body of the crossflow results in rotation. (b)  A gradient across the body of the body-parallel velocity does not result in rotation.
Figure 5: Rotational dynamics of the vortex ring model. (a)  A gradient along the body of the crossflow does not result in rotation. (b)  A gradient across the body of the body-parallel velocity results in rotation.
@@
III Passive dynamics of a pair
We present numerical examples of the three-dimensional dynamic interactions between a pair of identical swimmers. Such pairwise interactions can be thought of as the elemental interactions in a large school. To facilitate comparison with two-dimensional models, we focus on planar dynamics, where the swimmers’ bodies are co-planar initially and, therefore, for all time. This also allows us to fully characterize the phase space. We also include examples of non-planar dynamics. In what follows, lengths have been non-dimensionalized by the swimmer’s length $\ell$ and velocities have been non-dimensionalized by the self-propelled speed $U$.
@@ III.1 Symmetry reduction
With two swimmers, the dimension of the state space is 12, with six variables for each swimmer (three for the position of its center, and three for the unit normal describing its orientation). For the planar motion of two swimmers, however, leveraging symmetry allows us to describe the motion of the pair using only three variables. Specifically, due to a translational symmetry, only the separation between the swimmers’ centers (rather than their absolute positions) affects the dynamics. Similarly, a rotational symmetry makes it so that only the relative angle between the swimmers (rather than their absolute orientations) affects the dynamics. Without loss of generality, we place the swimmers in the $x$-$y$ plane, with one swimmer initially oriented in the positive $y$-direction, separated from each other by $(\Delta x,\Delta y)$, and with a relative angle $\Delta\theta$ between them, as in Figure 6. In this smaller three-dimensional symmetry-reduced phase space, we can fully characterize the possible dynamics. For non-planar dynamics, an additional angle is required to describe the dynamics.
@@
III.2 Planar dynamics of the three-dimensional model
@@
Figure 8: Planar trajectories of two swimmers for different initial conditions with $\Delta\theta=0$. (a)  $\Delta\theta=0$, $\Delta x=0.5$, $\Delta y=0$. In (a), the two swimmers’ bodies are represented by arrows whose colors represent time (darker is a later time). (b)  $\Delta\theta=0$, $\Delta x=5$, $\Delta y=0.5$. (c)  $\Delta\theta=0$, $\Delta x=1$, $\Delta y=1$. (d)  $\Delta\theta=0$, $\Delta x=0.5$, $\Delta y=1$.
@@
III.3 Planar dynamics of the two-dimensional model
@@
III.4 Non-planar dynamics of the three-dimensional model
For fully three-dimensional dynamics, an additional angle is required to describe all possible non-redundant initial configurations of a pair of swimmers. Without loss of generality, we place the center of the first swimmer at the origin and direct its unit vector $\mathbf{n}_{1}$ along the $x$-axis. The center of the second swimmer is placed in the plane $z=0$, and its unit vector $\mathbf{n}_{2}=R_{z}(\Delta\alpha)R_{y}(\Delta\beta)\mathbf{n}_{1}$ is produced by rotating $\mathbf{n}_{1}$ about the $y$-axis by $\Delta\beta$ and about the $z$-axis by $\Delta\alpha$. The rotation matrices are
$R_{z}(\Delta\alpha)=\begin{bmatrix}\cos\Delta\alpha&-\sin\Delta\alpha&0\\ \sin\Delta\alpha&\cos\Delta\alpha&0\\ 0&0&1\end{bmatrix},\quad R_{y}(\Delta\beta)=\begin{bmatrix}\cos\Delta\beta&0&\sin\Delta\beta\\ 0&1&0\\ -\sin\Delta\beta&0&\cos\Delta\beta\end{bmatrix}.$ (13)
On the basis of the planar results, we expect that the additional degree of freedom present in non-planar dynamics would increase the likelihood that the swimmers diverge since it allows for greater misalignment between the swimmers. An example of divergent motion is shown in Figure 14, where the swimmers’ paths have been projected onto three orthogonal planes for clarity. The swimmers initially interact with each other according to their mutually induced velocity fields before separating. Once far enough from each other, the swimmers move along approximately straight-line paths in directions differing from their initial orientations due to the initial interaction.
Figure 14: Non-planar trajectories of a pair of swimmers for different initial conditions. (a) $\Delta x=2$, $\Delta y=0$, $\Delta z=0$, $\Delta\alpha=-30$, $\Delta\beta=-15$. (b) $\Delta x=1.5$, $\Delta y=1$, $\Delta z=0$, $\Delta\alpha=-15$, $\Delta\beta=-15$. (c) $\Delta x=1$, $\Delta y=0.5$, $\Delta z=0$, $\Delta\alpha=-18$, $\Delta\beta=-18$.
Despite the potential for greater misalignment, we nevertheless observe configurations that lead to cohesive motions that are of the same character as in the planar case. In Figure 14 we show a three-dimensional braid-like motion analogous to the two-dimensional counterpart in Figure 8. While the planar braid consists of two scales of motion (the oscillation and the mean translation), the non-planar braid consists of three scales: the oscillation apparent in Figure 14, a mean straight-line motion, and a larger-scale oscillation that is approximately six times larger than the oscillation apparent in the plot. In Figure 14 we show a configuration that leads to more complex dynamics. For clarity, only a few oscillations are shown. Allowing the motion to evolve further in time reveals net translation of the pair of swimmers. This configuration leads to a three-dimensional complex braid-like motion analogous to the two-dimensional counterpart in Figure 8, although the non-planar complex braid again consists of three scales of motion (the oscillation apparent in Figure 14, a mean straight-line motion, and a larger-scale oscillation), whereas the planar complex braid consists of two scales of motion.
The non-planar dynamics are apparently very rich, and a detailed investigation will be the subject of future work. Most importantly, passively cohesive groups can exist in the fully three-dimensional problem.
III.5 Geometric effects
@@
Finally, we briefly study the effects of geometry by comparing long and thin swimmers (source-sink pairs) to short and fat swimmers (vortex ring model). The phase diagram for the planar dynamics of two initially aligned ($\Delta\theta=0$) vortex-ring swimmers is shown in Figure 15. The region where the vortex rings overlap is excluded. We observe four types of motions: in-line motion, divergence, collision, and oscillatory motion. Oscillatory motion is qualitatively similar to the braid-like motion in Figure 8, except that the paths of the swimmers do not intersect, instead following an alternating pattern of convergence and divergence. Although the types of motions are similar to what is observed for the source-sink model, the phase diagram differs substantially. First, the set of configurations leading to a cohesive group is much smaller than for the source-sink model. Second, cohesive configurations have small vertical separations, while for the source-sink model they had large vertical separations. Third, the vortex rings collide only when $\Delta y=0$; when they collide, they do so by turning head-on towards each other. These qualitative differences in the dynamics of short and fat swimmers compared to long and thin swimmers can be rationalized by appealing to how these two types of swimmers sample the velocity gradient, as was explained in Figures 4 and 5. Interestingly, the dynamics we observe in our vortex ring model are qualitatively similar to the dynamics observed for toroidal swimmers in Stokes flow [61], despite being at opposite ends of the Reynolds number spectrum.